# Random time-shift approximation enables hierarchical Bayesian inference of mechanistic within-host viral dynamics models on large datasets

### v2

@@ Abstract
Mechanistic mathematical models of within-host viral dynamics are tools for understanding how a virus’ biology and its interaction with the immune system shape the infectivity of a host. The biology of the process is encoded by the structure and parameters of the model that can be inferred statistically by fitting to viral load data. The main drawback of mechanistic models is that this inference is computationally expensive because the model must be repeatedly solved. This limits the size of the datasets that can be considered or the complexity of the models fitted. In this paper we develop a much cheaper inference method by implementing a novel approximation of the model dynamics that uses a combination of random and deterministic processes. This approximation also properly accounts for process noise early in the infection when cell and virion numbers are small, which is important for the viral dynamics but often overlooked. Our method runs on a consumer laptop and is fast enough to facilitate a full hierarchical Bayesian treatment of the problem with sharing of information to allow for individual level parameter differences. We apply our method to simulated datasets and a reanalysis of COVID-19 monitoring data in an National Basketball Association cohort of 163 individuals.
@@ 1 Introduction
Another way of looking at this approximation is that it allows fast forward simulation of the model without the need for costly stochastic simulation, for example using the Gillespie algorithm [22, 23]. For any given set of parameters, only a single solution of the deterministic model and samples from the corresponding time-shift distribution are required. In many existing studies of this problem the initial viral load is treated as being unobserved and hence a parameter to be inferred [11] or fixed at some value [5]. The time-shift idea facilitates a shift in perspective from inferring the initial variable viral load at a fixed time, to instead inferring a fixed viral load at a random time. This turns out to be a key insight as it allows the derivation of a number of further accurate approximations for the likelihood function that provide large computational speed ups, allowing inference on large datasets using a relatively complicated model.
@@ 2.2 Random time-shift approximation
with $(a,d,p)$ determining the particular member of the family. The quantity $\lambda$—the early growth rate of the model—and the time-shift parameters $\boldsymbol{\zeta}=(a,p,d)$ are numerically calculated from the branching process approximation to the CTMC (see [17] for concrete examples). The blue curve shown in Fig. 1 is this density evaluated for the parameters given in the caption. Recall that as $\tau$ is a function of $\boldsymbol{\theta}$, the parameters $\boldsymbol{\zeta}$ are also a function of $\boldsymbol{\theta}$ but that this has been suppressed in the notation for clarity. One final and important point is that the time-shift is independent of the infection time $t_{0}$.
@@ 2.5 Likelihood approximations
In this section, we outline the construction of the marginal likelihood for an individual, breaking it down into its components: the time-shift and the path likelihood—the likelihood function given a viral load (VL) sample path. This approach mirrors the general method used in the literature for constructing the likelihood in purely deterministic models [5, 11, 40].
@@
$h(\tau)=f(\boldsymbol{y}_{i}\,|\,\boldsymbol{z}_{i}(\tau),\boldsymbol{\theta}_{i},\kappa)f(\tau\,|\,\boldsymbol{\theta}_{i}).$ (9)
then the Laplace approximation to the marginal likelihood (Eq. (7)) is given by (see Section 4 of S1 text. for derivation)
$f(\boldsymbol{y}_{i}\,|\,\boldsymbol{\theta}_{i},\kappa)\approx\frac{\sqrt{2\pi}h(\tau_{0})}{\sqrt{-h^{\prime\prime}(\tau_{0})}},$ (10)
noting that $h(\tau)$ is assumed to be concave (implicit in applying the Laplace approximation), which implies $h^{\prime\prime}(\tau)<0$ around $\tau_{0}$. The quantity $\tau_{0}$, is the maximiser of Eq. (9) and can be estimated efficiently using optimisation. The denominator, $((-h^{\prime\prime}(\tau_{0}))^{-1/2}$, is the scale term and has a more complicated form but can also be computed through either numerical differentiation or automatic differentiation of Eq. (9). In this work, all our code is written in the Julia programming language [42], and the package ForwardDiff.jl [43] enables efficient computation of exact derivatives through forward mode automatic differentiation. In Section 4 of S1 text., we plot profile likelihoods—i.e., the likelihood with all parameters fixed except for one, which is allowed to vary—for a simulated VL trajectory to show the accuracy of the approximation in Eq. (10) versus the exact calculation in Eq. (7). The profile likelihoods had a maximum (absolute) error of $10^{-7}$ at the tested points, which indicates strong agreement between our Laplace approximation and the exact likelihood.
@@ 2.6 Priors
In this section, we detail our choices of priors and fixed parameters. We center our priors at values consistent with studies of other viral infections (e.g., influenza) and COVID-19, and use weakly informative priors with higher variance to account for uncertainty in these estimates. Prior predictive checks confirm that the effective priors on individual parameters remain weakly informative. Our primary focus in this work is methodological—demonstrating how to incorporate stochasticity into WHVD models and speed up the inference—rather than conducting a thorough analysis of this data. Moreover, because our hierarchical model is fit to a larger number of individuals than in previous studies, the population-level estimates are better informed by the data and less reliant on the prior. While hierarchical models can be sensitive to prior misspecification, the larger sample size here mitigates this concern, as the data dominate the inference.
Due to identifiability concerns, some VL parameters are fixed at values that are well established in the literature. We assume that the eclipse phase of newly infected cells, $k_{i}$, and the clearance rate of virus, $c_{i}$, are fixed at $k_{i}=4\textrm{ d}^{-1}$ and $c_{i}=10\textrm{ d}^{-1}$ for $i=1,\dots,N$, respectively [5]. The individual-level parameters, $R_{0,i}$, $\delta_{i}$, and $\rho_{i}$, are all assumed to be normally distributed (truncated to be non-negative) with unknown mean and scale parameters. Here we detail the hyper-priors on these parameters. As outlined in Section 2.2, we parameterise the model in terms of the basic reproduction number, $R_{0}$. Table 2 gives some estimates of $R_{0}$ for COVID-19 and Influenza as well as the dataset size (number of individuals) from the literature. None of these studies implemented hierarchical models and instead fit each individual time-series.
Estimate (95% CI) Number of individuals Virus Reference Dataset
8.2 (NA) 25 COVID-19 [5] NBA
11.1 (6.6, 18.5) 6 Influenza A [2] Other
13.51 (4.5, 45.12) 9 COVID-19 [44] Other
14.2 (NA) 6 COVID-19 [18] NBA
Table 2: Estimated $R_{0}$ values from published literature. Estimates and 95% CI where available of $R_{0}$ from published models. Information as to the type of viral infection and the number of individuals in the study are also provided. The dataset column indicates what dataset was of interest for that particular study, with NBA denoting a subset of our dataset and other denoting a different dataset.
@@
This choice results in a mean of 10 with a 95% quantile interval of $(2.3,23.3)$, providing a weakly informative prior centered near commonly reported estimates.
A meta-analysis of WHVD studies on Influenza [45] provides experimental estimates for the average period of infectiousness of an infected cell, $1/\delta$, ranges from 0.5 to 2 days which corresponds to a removal rate of between $0.5$ and $2$. We assume
$\mu_{\delta}\sim\textrm{Gamma}(26,0.05),$
which yields a mean rate of removal of $1.3~d^{-1}$ with a 95% quantile interval of $(0.85,1.85)$. The parameter $\mu_{\rho}$ is difficult to assign a prior, as it has identifiability issues similar to $\beta$ [5, 18]. To address this, we assume
$\mu_{\rho}\sim\textrm{Gamma}(10,0.3),$
which has median $2.9~d^{-1}$ (roughly based on [5]) and a 95% quantile interval of $(1.44,5.13)$. This distribution is deliberately weakly informative and reflects values estimated for COVID-19 in other studies [46, 44].
@@ 2.8 Datasets
In this analysis, we use a collection of datasets: a published dataset of COVID-19 VL measurements collected during the 2020-2021 NBA season restart, and 50 synthetic datasets for method validation. Each synthetic dataset serves as a controlled setting with frequent, low-noise observations and known parameters. The purpose of the simulated datasets is to validate the method can recover parameters in an ideal setting. Each dataset consists of $N=100$ individual timeseries and are generated using our prior model for the hyperparameters $\boldsymbol{\phi}$. The full data generation process is detailed in Section 2.8.1, and an example simulated cohort is shown in Fig. 2A.
@@
The limit of detection (LOD) in the NBA dataset is 2.658 log copies/mL and frequent testing and individuals experiencing only single infections resulted in many observations (i.e. some individuals having 80% LOD observations) at this threshold for each individual. Since it is most likely that individuals are only actually positive for some number of days pre/post the time where they have above LOD observations, we truncated the time series, retaining only the LOD measurements such that above-threshold observations remained between them. In some cases, multiple values below the LOD occur before another measurement slightly above the detection limit which complicates fitting. To address this, we further truncated time series at three consecutive threshold observations. Finally, there were some short time-series—i.e. less than 2 observations above the LOD—and these were excluded. After filtering our dataset contains 163 of the original 241 timeseries. Note that we use the same post-processing procedure and LOD for the synthetic dataset. The NBA data is shown in Fig. 2B alongside an example simulated cohort in panel Fig. 2A. The NBA data (Fig. 2B) exhibits visibly higher variance across the entire time horizon compared to the simulated dataset (Fig. 2A). This is primarily because the synthetic data were deliberately generated with lower noise to enable more controlled validation of our method.
@@
Our hierarchical model shares information from individuals with high information to improve estimates for those with sparse data, a process known as partial pooling [51, Chapter 5]. Previous analyses of this dataset [5, 12, 4, 6] used deterministic models [5, 12] or piecewise linear functions [12, 4, 6], whereas our model explicitly accounts for stochasticity in VL trajectories. Since we include individuals with both high and low information, direct parameter comparisons to the model presented in the Supplementary Information (SI) of Zitzmann et al. [5] are not possible. However, we can compare our $R_{0}$ estimates to their results and those from other SARS-CoV-2 studies [52] as well as other respiratory illnesses [31, 52]. Additionally, we compare predicted VL trajectories to those of some of the SI fits in Zitzmann et al. [5]. Phenomenological models [11, 4] estimate parameters based on piecewise linear models, which are not directly comparable to our mechanistic approach, but our model can estimate the same summary statistics as phenomenological models while also inferring biological quantities like $R_{0}$.
@@
Fig 2: Scatter plots of an example simulated dataset and the NBA dataset. Panel (a): an example synthetic VL dataset generated according to the process in Section 2.8.1. Panel (b): the NBA VL dataset for the 2020-2021 seasons. In each panel an individual data point corresponds to a single measurement for a single individual. Points are shown as semi-transparent to approximate the “density” of points—highlighting outliers in panel (b). The red dashed line indicates the limit of detection.
@@ 2.8.1 Generating synthetic datasets
We construct our synthetic datasets by initially sampling VL model parameters from the prior setup as outlined in Section 2.6. This requires fixing the shared parameters and sampling individual parameters from priors. The means of the shared parameters ($\mu_{R_{0}},\mu_{\delta},\mu_{\rho}$) are chosen to be in agreement with previous estimates [5, SI]. The scale parameters ($\sigma_{R_{0}},\sigma_{\delta},\sigma_{\rho}$) were arbitrarily selected based on comparing the simulations to the NBA dataset. The measurement noise parameter $\kappa$ was chosen to be $0.5$ to reflect a reasonable amount of noise in the measured VLs but still a “low noise” situation to assess the model in an ideal scenario. The values of all shared parameters are given in Table 3.
@@
Model parameters $\boldsymbol{\theta}_{i}$ are sampled for each individual from the priors assuming the values of the hyper-parameters in Table 3. Infection times are then drawn from the prior and then a time-series for each individual is simulated by running Algorithm 1. The variability in the priors causes a wide range of growth and decline rates as well as peak log-VL values which can cause unrealistic scenarios such as virus never being eliminated. In order to ensure we generate realistic data, we simulate under a rejection framework whereby we accept a simulation if an individual has peak log-VL between $5$ and $10$ (visually consistent with the range in the peak log-VLs in the NBA dataset) and a corresponding peak infection time of between 4 and 9 days post-infection. We also require that the decline rate is consistent with a typical acute infection and so we require the log-VL to be less than $4.5$ at the end of the observation period (21 days post-infection). This condition ensures we do not have long-tails which are biologically invalid (i.e. no viral clearence). As in the NBA data, we assume a LOD of 2.658 log copies/mL, processing these as outlined in Section 2.8. Overall, the simulated datasets constitute an idealised scenario where individuals have multiple observations in both the growth and decline phases.
@@ 3 Results
In this section we apply our method to the datasets outlined in Section 2.8. The initial exploration of the model fit to the synthetic datasets enables us to demonstrate the ability of the algorithm to produce accurate fits to the data in a situation where we know the true parameter values. The NBA dataset is then used to demonstrate the utility of the method in a real-world scenario where the true parameter values are unknown.
All simulations and models were implemented in Julia [42] and run on a 2021 MacBook Pro with an M1 chip and 16 GB of RAM. Likelihood computations were parallelised across 4 high-performance CPU cores, as described in Section 5 of S1 text.. Pilot runs were used to tune the proposal distributions for both individual and shared parameters. Following this, we initialised three chains from perturbed versions of the parameters used to generate the synthetic data, to assess the robustness of the sampler to initial conditions. Each of the chains were run for 100,000 iterations, with the first 20,000 discarded as burn-in. The run times for the NBA dataset (the larger of the datasets) took approximately 10 minutes per chain for a total runtime of approximately 30 minutes.
Trace plots for all shared parameters and a subset of individual parameters were visually inspected for divergences and showed no signs of poor mixing across the chains for both analyses. All parameters had $\hat{R}\leq 1.01$, indicating good within- and between-chain mixing [53], and minimum effective sample sizes exceeding 500. The acceptance rates were approximately $0.30$ for the individual (averaged across groups) and shared parameters, respectively, both within the expected range for well-tuned random walk Metropolis proposals [54].

### v3

@@ Abstract
Mechanistic mathematical models of within-host viral dynamics are tools for understanding how a virus’ biology and its interaction with the immune system shape the infectivity of a host. The biology of the process is encoded by the structure and parameters of the model that can be inferred statistically by fitting to viral load data. The main drawback of mechanistic models is that this inference is computationally expensive because the model must be repeatedly solved. This limits the size of the datasets that can be considered or the complexity of the models fitted. In this paper we develop a much cheaper inference method for this class of models by implementing a novel approximation of the model dynamics that uses a combination of random and deterministic processes. This approximation also properly accounts for process noise early in the infection when cell and virion numbers are small, which is important for the viral dynamics but often overlooked. Our method runs on a consumer laptop and is fast enough to facilitate a full hierarchical Bayesian treatment of the problem with sharing of information to allow for individual level parameter differences. We apply our method to simulated datasets and a reanalysis of COVID-19 monitoring data in an National Basketball Association cohort of 163 individuals.
@@ 1 Introduction
Another way of looking at this approximation is that it allows fast forward simulation of the model without the need for costly stochastic simulation, for example using the Gillespie algorithm [22, 23]. For any given set of parameters, only a single solution of the deterministic model and samples from the corresponding time-shift distribution are required. In many existing studies of this problem the initial viral load is treated as being unobserved and hence a parameter to be inferred [11] or fixed at some value [5]. The time-shift idea facilitates a shift in perspective from inferring the initial variable viral load at a fixed time, to instead inferring a fixed viral load at a random time. This turns out to be a key insight as it allows the derivation of a number of further accurate approximations for the likelihood function that provide large computational speed ups over standard stochastic simulation techniques, allowing inference on large datasets using a relatively complicated model.
@@ 2.2 Random time-shift approximation
with $(a,d,p)$ determining the particular member of the family. The quantity $\lambda$, representing the early growth rate of the model, and the time-shift parameters $\boldsymbol{\zeta}=(a,p,d)$ are numerically calculated from the branching process approximation to the CTMC (see [17] for concrete examples). The blue curve shown in Fig. 1 is this density evaluated for the parameters given in the caption. Recall that as $\tau$ is a function of $\boldsymbol{\theta}$, the parameters $\boldsymbol{\zeta}$ are also a function of $\boldsymbol{\theta}$ but that this has been suppressed in the notation for clarity. One final and important point is that the time-shift is independent of the infection time $t_{0}$.
@@ 2.3 Viral load measurement model
Note that predicted viral loads below one particle (i.e. $V(t)<1$) were mapped to a small pseudocount ($10^{-5}$) before log-transformation to avoid numerical instabilities.
@@ 2.5 Likelihood approximations
In this section, we outline the construction of the marginal likelihood for an individual, breaking it down into its components: the time-shift and the path likelihood, the likelihood function given a viral load (VL) sample path. This approach mirrors the general method used in the literature for constructing the likelihood in purely deterministic models [5, 11, 40].
@@
$h(\tau)$ $=f(\boldsymbol{y}_{i}\,|\,\boldsymbol{z}_{i}(\tau),\boldsymbol{\theta}_{i},\kappa)f(\tau\,|\,\boldsymbol{\theta}_{i}),$ (9)
and let $g(\tau)=\log{h(\tau)}$. Then the Laplace approximation to the marginal likelihood (Eq. (7)) is given by (see Section 4 of S1 text. for derivation)
$f(\boldsymbol{y}_{i}\,|\,\boldsymbol{\theta}_{i},\kappa)\approx\frac{\sqrt{2\pi}h(\tau_{0})}{\sqrt{-g^{\prime\prime}(\tau_{0})}},$ (10)
noting that $g(\tau)$ is assumed to be concave (implicit in applying the Laplace approximation), which implies $g^{\prime\prime}(\tau)<0$ around $\tau_{0}$. The quantity $\tau_{0}$, is the maximiser of Eq. (9) and can be estimated efficiently using optimisation. The denominator, $(-g^{\prime\prime}(\tau_{0}))^{-1/2}$, is the scale term and has a more complicated form but can also be computed through either numerical differentiation or automatic differentiation of the log of Eq. (9). In this work, all our code is written in the Julia programming language [42], and the package ForwardDiff.jl [43] enables efficient computation of exact derivatives through forward mode automatic differentiation. In Section 4 of S1 text., we plot profile likelihoods—i.e., the likelihood with all parameters fixed except for one, which is allowed to vary—for a simulated VL trajectory to show the accuracy of the approximation in Eq. (10) versus the exact calculation in Eq. (7). The profile likelihoods had a maximum (absolute) error of $10^{-7}$ at the tested points, which indicates strong agreement between our Laplace approximation and the exact likelihood.
@@ 2.6 Priors
In this section, we detail our choices of priors and fixed parameters. We center our priors at values consistent with other studies of COVID-19, and use weakly informative priors with higher variance to account for uncertainty in these estimates. Prior predictive checks confirm that the effective priors on individual parameters remain weakly informative. Our primary focus in this work is methodological, demonstrating how to incorporate stochasticity into WHVD models and speed up the inference, rather than conducting a thorough analysis of this data. Moreover, because our hierarchical model is fit to a larger number of individuals than in previous studies, the population-level estimates are better informed by the data and less reliant on the prior. While hierarchical models can be sensitive to prior misspecification, the larger sample size here mitigates this concern, as the data dominate the inference.
Due to identifiability concerns, some VL parameters are fixed at values that are well established in the literature. We assume that the eclipse phase of newly infected cells, $k_{i}$, and the clearance rate of virus, $c_{i}$, are fixed at $k_{i}=4\textrm{ d}^{-1}$ and $c_{i}=10\textrm{ d}^{-1}$ for $i=1,\dots,N$, respectively [5]. The individual-level parameters, $R_{0,i}$, $\delta_{i}$, and $\rho_{i}$, are all assumed to be normally distributed (truncated to be non-negative) with unknown mean and scale parameters. Here we detail the hyper-priors on these parameters. As outlined in Section 2.2, we parameterise the model in terms of the basic reproduction number, $R_{0}$. Table 2 gives some estimates of $R_{0}$ for COVID-19 and the dataset size (number of individuals) from the literature. None of these studies implemented hierarchical models and instead fit each individual time-series.
Estimate (95% CI) Number of individuals Reference Dataset
8.2 (NA) 25 [5] NBA
13.51 (4.5, 45.12) 9 [44] Other
14.2 (NA) 6 [18] NBA
Table 2: Estimated $R_{0}$ values from published literature. Estimates and 95% CI where available of $R_{0}$ from published models. The number of individuals in the study are also provided. The dataset column indicates what dataset was of interest for that particular study, with NBA denoting a subset of our dataset and other denoting a different dataset.
@@
This choice results in a mean of 10 with a 95% quantile interval of $(2.3,23.3)$, providing a weakly informative prior centered near the commonly reported estimates.
To parameterise the priors on the mean infectious period, $\mu_{\delta}$, and the mean viral replication rate, $\mu_{\rho}$, we draw on previous studies fitting similar models to COVID-19 data (see [46, 44] and the SI of [5]). We assume
$\mu_{\delta}\sim\text{Gamma}(26,0.05),$
which corresponds to a mean removal rate of $1.3~\mathrm{d^{-1}}$ and a 95% interval of $(0.85,1.85)$, consistent with published estimates. The parameter $\mu_{\rho}$ is difficult to assign a prior due to identifiability issues similar to those faced by $\beta$ [5, 18]. To address this, we assume
$\mu_{\rho}\sim\text{Gamma}(10,0.3),$
with median $2.9~\mathrm{d^{-1}}$ (roughly based on fits in the SI of [5]) and a 95% interval of $(1.44,5.13)$. This choice reflects prior knowledge while remaining flexible enough for the data to inform the estimate. Note that the main-text model in [5] includes an immune response, leading to a much larger magnitude ($\sim 100$), so direct numerical comparison is not appropriate.
@@ 2.8 Datasets
In this analysis, we use a collection of datasets: a published dataset of COVID-19 VL measurements collected during the 2020-2021 NBA season restart, and 200 synthetic datasets for method validation. Each synthetic dataset serves as a controlled setting with frequent, low-noise observations and known parameters. The purpose of the simulated datasets is to validate the method can recover parameters in an ideal setting. Each dataset consists of $N=100$ individual timeseries and are generated using our prior model for the hyperparameters $\boldsymbol{\phi}$. The full data generation process is detailed in Section 2.8.1, and an example simulated cohort is shown in Fig. 2A.
@@
The limit of detection (LOD) in the NBA dataset is 2.658 log copies/mL and frequent testing and individuals experiencing only single infections resulted in many observations (i.e. some individuals having 80% LOD observations) at this threshold for each individual. Since it is most likely that individuals are only actually positive for some number of days pre/post the time where they have above LOD observations, we truncated the time series, retaining only the LOD measurements such that above-threshold observations remained between them. In some cases, multiple values below the LOD occur before another measurement slightly above the detection limit which complicates fitting. To address this, we further truncated time series at three consecutive threshold observations. Finally, there were some short time-series with less than 2 observations above the LOD and these were excluded. After filtering our dataset contains 163 of the original 241 timeseries. Note that we use the same post-processing procedure and LOD for the synthetic dataset. The NBA data is shown in Fig. 2B alongside an example simulated cohort in panel Fig. 2A. The NBA data (Fig. 2B) exhibits visibly higher variance across the entire time horizon compared to the simulated dataset (Fig. 2A). This is primarily because the synthetic data were deliberately generated with lower noise to enable more controlled validation of our method.
@@
Our hierarchical model shares information from individuals with high information to improve estimates for those with sparse data, a process known as partial pooling [51, Chapter 5]. Previous analyses of this dataset [5, 12, 4, 6] used deterministic models [5, 12] or piecewise linear functions [12, 4, 6], whereas our model explicitly accounts for stochasticity in VL trajectories. Since we include individuals with both high and low information, direct parameter comparisons to the model presented in the SI of Zitzmann et al. [5] are not possible. However, we can compare our $R_{0}$ estimates to their results and those from other SARS-CoV-2 studies [52] as well as other respiratory illnesses [31, 52]. Additionally, we compare predicted VL trajectories to those of some of the SI fits in Zitzmann et al. [5]. Phenomenological models [11, 4] estimate parameters based on piecewise linear models, which are not directly comparable to our mechanistic approach, but our model can estimate the same summary statistics as phenomenological models while also inferring biological quantities like $R_{0}$.
@@
Fig 2: Scatter plots of an example simulated dataset and the NBA dataset. (A) An example synthetic VL dataset generated according to the process in Section 2.8.1. (B) The NBA VL dataset for the 2020-2021 seasons. In each panel an individual data point corresponds to a single measurement for a single individual. Points are shown as semi-transparent to approximate the “density” of points, highlighting outliers in panel (B). The red dashed line indicates the limit of detection.
@@ 2.8.1 Generating synthetic datasets
We generated synthetic viral load (VL) datasets by sampling model parameters from the priors described in Section 2.6, fixing the shared (hyper-)parameters and drawing individual-level parameters accordingly (e.g. for the basic reproduction numbers, $R_{0,i}\sim\mathcal{N}(\mu_{R_{0}},\sigma_{R_{0}})$ and similarly for $\delta_{i}$ and $\rho_{i}$). The shared parameter means ($\mu_{R_{0}}$, $\mu_{\delta}$, $\mu_{\rho}$) were chosen to align with previous estimates [5, SI], while their corresponding scales ($\sigma_{R_{0}}$, $\sigma_{\delta}$, $\sigma_{\rho}$) were selected heuristically based on comparisons between simulated and NBA datasets. The measurement noise parameter $\kappa$ was set to $0.5$, representing a moderate, “low-noise” scenario suitable for evaluating model performance under idealised conditions. The full set of hyper-parameter values is given in Table 3.
@@
For each individual, infection times were drawn from the prior, and corresponding time series were simulated using Algorithm 1. Because variability in the priors can produce unrealistic trajectories (e.g., non-clearing infections or implausible peak VLs), we applied a rejection-sampling procedure to retain only biologically plausible cases. Specifically, we accepted simulations with peak log-VL between 5 and 10 (consistent with the NBA dataset), peak time between 4 and 9 days post-infection, and log-VL below 4.5 at 21 days post-infection to ensure viral clearance. As in the NBA data, a limit of detection (LOD) of 2.658 log copies/mL was imposed and processed as described in Section 2.8. The resulting datasets thus represent an idealised setting in which individuals have well-resolved observations across both the growth and decline phases of infection.
@@ 3 Results
In this section, we apply our method to the datasets described in Section 2.8. We first analyse the synthetic datasets to demonstrate the algorithm’s ability to recover accurate parameter estimates in a setting where the true values are known. We then apply the method to the NBA dataset to illustrate its utility in a real-world scenario where the true parameter values are unknown.
All simulations and models were implemented in Julia v1.12 [42] and executed on a 2021 MacBook Pro with an M1 chip and 16 GB of RAM. Likelihood computations were parallelised across four high-performance CPU cores, as described in Section 5 of S1 text.. Pilot runs were used to tune the proposal distributions for both individual and shared parameters. Following the pilot runs, we initialised three chains from perturbed versions of the posterior means obtained in the pilot phase. For each simulated dataset, each chain was run for 100,000 iterations, with the first 20,000 discarded as burn-in. For the NBA dataset, each chain was run for 200,000 iterations, discarding the first 20,000 as burn-in. The runs for the NBA dataset (the larger of the two) required approximately 15 minutes per chain, giving a total runtime of about 45 minutes.
Trace plots for all shared parameters and a subset of individual parameters were visually inspected for divergences and showed no evidence of poor mixing across chains in either analysis. All parameters had $\hat{R}\leq 1.01$, indicating good within- and between-chain convergence [53], with minimum effective sample sizes exceeding 500. The acceptance rates were approximately 0.30 for both the individual (averaged across groups) and shared parameters, within the expected range for well-tuned random walk Metropolis proposals [55].